• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Holmstrom

    Log entry time 00:08:34 on July26,2003

    Entry number 106903

    keyword=Swing Shift Summary

    16:00 Start Swing Shift -- No Beam, we are in Controlled Access, no Hall A Main control. MCC is working on a problem with the fast feedback.

    17:30 We did an EPR measurement, with P=38.34%.

    17:45 We did an NMR measurement, with P=39.1%.

    18:20 Hall A Main control back.

    18:45 MCC called and they want to send us Beam for tunning.

    19:05 MCC called and Asked us to turn off the Septum and Q1 on the right HRS.

    19:40 Pulsed Beam into the Hall.

    20:17 MCC called and they will do a controled access to remove the Beam Dump Ion Chamber.

    20:19 Began ramping the Septum and Q1 back up.

    20:52 Jing changed the voltages on the right arm Shower/Preshower and Gas Charenkov, in preperation for a detector check.

    21:05 Beam Permit.

    21:10 Beam Returns.

    21:15 Beam Dump studies are over. MCC says they will work on the Beam size.

    22:30 Beam Gone

    23:10 Compton Tune
    23:20 Moller Starts Tune

    23:40 Talk to J.P., I messed up the Harp Scan reading. We need to spread in Y and narrow in X.